OpenStudy (anonymous):

Solve each equation for 0

OpenStudy (campbell_st):

well remove the common factor and you have cos(x) (cos(x) -1) = 0 so you need to find where cos(x) = 0 and cos(x) -1= 0 hope it helps

OpenStudy (anonymous):

what is the common factor?

OpenStudy (anonymous):

i mean in this equation\

OpenStudy (campbell_st):

its cos(x) cos(x)(cos(x) - 1) = cos^2(x) - cos(x)